# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@ 6. Challenges and Future Directions
The increasing popularity of foundation models on vision and language tasks has also given rise to the emergence of various multimodal foundation models(Radford et al., 2021). While many of these models demonstrate strong FSL performance or have been adapted to be effective in that setting, their shortcomings in achieving holistic and human-like generalization capabilities are still notable. Broadly speaking, generalization can be multi-faceted and may be abstracted in different levels based on context:
  1. (1)
@@
    Single instance of a learning problem. Better known as domain adaptation/generalization, the model’s goal is to be able to generalize strongly to data from a previously unseen distribution for a specific problem.
  2. (2)
@@
While the first two cases have been extensively researched for FSL, only recently has generalization within a specific modality been explored via models that can generalize across different dense visual prediction tasks(Kim et al., 2023), and the curation of FSL datasets containing multiple visual learning problems (Bohdal et al., 2023). However, current vision learning models struggle to generalize to new learning problems, and overall, FSL methods still face challenges in extrapolating to unseen modalities. Creating a genuine generalist model capable of FSL across novel modalities could be an ambitious but exciting future direction.
Human-like learning
@@
One of the major attributes of human intelligence is the ability to learn new concepts quickly, master new tasks, and combine them to solve complex problems, given only a few examples, possibly due to humans’ inherent inductive biases and their ability to leverage multimodal signals(Lin et al., 2023). However, current machine learning models struggle with fast adaptation to complex tasks and lack human-like intelligence attributes such as compositionality, concept association, relational understanding, and conceptual meaning(Jiang et al., 2023a). Systematicity also remains a challenge for modern models(Lake and Baroni, 2023).
@@
Rapid adaptation is generally prevalent in nature and has been studied for many decades in cognitive science, psychology, and neuroscience, with results highlighting the existence of different and nested timescales of learning. These observations have also been utilized in meta-learning via hierarchical Bayesian models(Lake et al., 2015). The compositionality and expressiveness of natural language have also been leveraged for FSL of concepts, focusing on induction and utilizing a human-like prior extracted from behavioral data from humans(Ellis, 2024). Moreover, meta-learning has been used for human-like learning of algebraic systems, leading to models that learn how to generalize and adopt human inductive biases(Lake and Baroni, 2023) systematically. On a different direction, brain mechanisms such as synapse formation in the human brain have inspired novel approaches in FSL, such as adaptation mechanisms for few-shot generative domain adaptation, enhancing the model with human-like properties such as memory and domain association(Wu et al., 2023b). Lastly, it has been shown that increased AI alignment can improve FSL models, adaptation to new domains, and robustness to adversarial examples, possibly by serving as a source of inductive bias that reduces the amount of information needed from data(Sucholutsky and Griffiths, 2024).
@@
It is worth noting that while cognitive science and neuroscience have focused on discovering the mechanisms of human intelligence, machine learning focuses on replicating these capabilities. In that direction, a new multimodal dataset for FSL has been proposed that enables few-shot multimodal abstract reasoning evaluation of models over various cognitive tasks(Jiang et al., 2023a). Preliminary results on the dataset suggest significant misalignment between machines and humans regarding FSL capabilities, raising questions on how to align machine and human learning and create agents with human-like FSL capabilities. Consequently, methods combining neuroscience and cognitive science with machine learning might be the way forward.
7. Conclusion
@@
Although deep learning has traditionally relied on large amounts of data to facilitate learning, there is an increasing focus on techniques that can learn with limited data. Similarly, despite the trend of creating large foundation models trained on massive datasets, there is a growing interest in exploring new few-shot learning (FSL) techniques and applying them in practical settings. Given FSL’s recent rapid expansion, this paper aims to provide an overview of this research field, covering its various aspects. The first step is to frame FSL by providing a formal definition based on which different proposed methods are developed. To position FSL within the broader machine learning paradigm, its connection with different learning fields is also examined. At the same time, models from older and more recent learning paradigms are presented to better understand their interactions and the range of approaches available based on a novel taxonomy. FSL’s bloom, which is also reflected in the numerous real-world applications in which it has been applied, is also described. Additionally, various emerging trends are discussed, including in-context learning and the rise of new applications in previously unexplored areas. As a final step, we aim to encourage further research by identifying key areas where FSL can have a significant impact, such as Green AI and bridging the gap between machine learning and human intelligence.
Acknowledgements.

### v3

@@ 6.7 Fully generalizable FSL
The increasing popularity of foundation models on vision and language tasks has also given rise to various multimodal foundation models [167]. While many of these models demonstrate strong FSL performance or have been adapted to be effective in that setting, their shortcomings in achieving holistic and human-like generalization capabilities are still notable. Broadly speaking, generalization can be multi-faceted and may be abstracted in different levels based on context:
  1. 1.
@@
    Single instance of a learning problem. Better known as domain adaptation/generalization, the model’s goal is to generalize strongly to data from a previously unseen distribution for a specific problem.
  2. 2.
@@
While the first two cases have been extensively researched for FSL, only recently has generalization within a specific modality been explored via models that can generalize across different visual prediction tasks [107, 10], and the curation of FSL datasets containing multiple visual learning problems [18]. However, current vision learning models struggle to generalize to new learning problems, and overall, FSL methods still face challenges in extrapolating to unseen modalities. Creating a genuine generalist model capable of FSL across novel modalities could be an ambitious but exciting future direction.
6.8 Human-like learning
@@
One of the major attributes of human intelligence is the ability to learn new concepts quickly, master new tasks, and combine them to solve complex problems, given only a few examples, possibly due to humans’ inherent inductive biases and their ability to leverage multimodal signals [130]. However, current machine learning models struggle with fast adaptation to complex tasks and lack human-like intelligence attributes such as compositionality, concept association, relational understanding, and conceptual meaning [92]. Systematicity also remains a challenge for modern models [115].
@@
Rapid adaptation is generally prevalent in nature and has been studied for many decades in cognitive science, psychology, and neuroscience, with results highlighting the existence of different and nested timescales of learning. These observations have also been utilized in meta-learning via hierarchical Bayesian models [116]. The compositionality and expressiveness of natural language have also been leveraged for FSL of concepts, focusing on induction and utilizing a human-like prior extracted from behavioral data from humans [49]. Moreover, meta-learning has been used for human-like learning of algebraic systems, leading to models that learn how to generalize and adopt human inductive biases [115] systematically. On a different direction, brain mechanisms such as synapse formation in the human brain have inspired novel approaches in FSL, such as adaptation mechanisms for few-shot generative domain adaptation, enhancing the model with human-like properties such as memory and domain association [229]. Lastly, it has been shown that increased AI alignment can improve FSL models, adaptation to new domains, and robustness to adversarial examples, possibly by serving as a source of inductive bias that reduces the amount of information needed from data [191].
@@
It is worth noting that while cognitive science and neuroscience have focused on discovering the mechanisms of human intelligence, machine learning focuses on replicating these capabilities. In that direction, a new multimodal dataset for FSL has been proposed that enables few-shot multimodal abstract reasoning evaluation of models over various cognitive tasks [92]. Preliminary results on the dataset suggest significant misalignment between machines and humans regarding FSL capabilities, raising questions on how to align machine and human learning and create agents with human-like FSL capabilities. Consequently, methods combining neuroscience and cognitive science with machine learning might be the way forward.
7 Conclusion
@@
Although deep learning has traditionally relied on large amounts of data to facilitate learning, there is an increasing focus on techniques that can learn with limited data. Similarly, despite the trend of creating large foundation models trained on massive datasets, there is a growing interest in exploring new FSL techniques and applying them in practical settings. Given FSL’s recent rapid expansion, this paper aims to provide an overview of this research field, covering its various aspects. The standard FSL formulation is initially provided, based on which different proposed methods are developed. To position FSL within the broader machine learning paradigm, its connection with different learning fields is also examined. Subsequently, we present a comprehensive taxonomy of FSL methods that extends existing taxonomies with recent approaches and model families and also introduces emerging learning paradigms within FSL, such as in-context learning. Additionally, we provide a broad overview of hybrid approaches that extend FSL beyond the presented supervised learning setting. FSL’s bloom, which is also reflected in the numerous real-world applications in which it has been applied, is also described. Finally, we discuss various emerging trends, including the few-shot evaluation of foundation models and the application of meta-learning in novel settings, and aim to encourage further research by identifying key areas where FSL can have a significant impact, such as Green AI and bridging the gap between machine learning and human intelligence.
[…]
Acknowledgements